OpenStudy (anonymous):

An art-supply store has a sale on oil paint. Some paints are $6.00 per tube and some paints are $4.50 per tube. An artist spends $96 on 18 tubes. How many of each type were purchased?

OpenStudy (anonymous):

so its not A or C that what I got so far , can someone help me with the answer

OpenStudy (anonymous):

here are the choices A) 8 tubes at $ 6.00 each; 8 tubes at $ 4.50 each B) 8 tubes at $ 6.00 each; 10 tubes at $ 4.50 each C) 18 tubes at $ 6.00 each; 8 tubes at $ 4.50 each D) 10 tubes at $ 6.00 each; 8 tubes at $ 4.50 each

OpenStudy (anonymous):

D

OpenStudy (anonymous):

10*6=60 8*4.50=36 60+36=96

OpenStudy (anonymous):

got it thanks @myx